Determine the value of a, so that If α=2i+aj+k and \eta=4i−j−2k are perpendicular.

a.2
b.5
c.4
d.3
1
Expert's answer
2020-03-20T17:48:57-0400

Vectors "\alpha=(2,a,1)" and "\beta = (4,-1,-2)" are perpendicular if and only if their dot product is equal to zero.


"(2,a,1)\cdot(4,-1,-2)=0""2\cdot 4+a\cdot(-1)+1\cdot(-2)=0""8-a-2=0""6-a=0""a=6"

Answer: 6